OpenStudy (anonymous):

can someone show me how to convert 500 cm/ms to m/s please and how to convert in steps thank you

OpenStudy (calculusxy):

What do you mean by "cm/ms to m/s"?

OpenStudy (austinl):

\(\dfrac{500cm}{1ms} \times \dfrac{1meter}{100cm} \times \dfrac{1000ms}{1sec} = Your~Answer\)

OpenStudy (anonymous):

Thank you

OpenStudy (austinl):

No problem, and that is what is known as dimensional analysis.